### v2

# Correlators are simpler than wavefunctions

Abstract
Several recent works have revealed a simplicity in equal-time correlators that is absent in their wavefunction counterparts. In this letter, we show that this arises from the simple fact that the correlator is obtained by integrating Feynman propagators over the full spacetime, as opposed to the half-space for the wavefunction. Several striking new properties of correlators for any graph are made obvious from this picture. Certain patterns of poles that appear in the wavefunction do not appear in the correlator. The correlator also enjoys several remarkable factorization properties in various limits. Most strikingly, the correlator admits a systematic Laurent expansion in the neighborhood of every pole, with the first subleading term vanishing for every pole. There is an especially simple understanding of the expansion around the total energy pole up to second order, given by a differential operator acting on the amplitude. Finally, we show how these results extend beyond single graphs to the full correlator in Tr $\phi^{3}$ theory.
I Introduction
The flat-space vacuum wavefunction for scalar theories with polynomial interactions has attracted considerable interest in recent years [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36], due to its close connection with the cosmological wavefunction in simple toy model cosmologies. Defined by the path integral [10]:
$\Psi[\Phi]$ $=\int_{\varphi(t=-\infty)}^{\varphi(t=0)}\mathcal{D}\varphi\>e^{-i\int_{-\infty}^{0}dt\,\int_{-\infty}^{+\infty}d^{n}x\,S[\varphi]},$ (1)
the wavefunction encodes the quantum state of the system on the $t=0$ slice. The physically relevant observables, however, are the equal-time correlation functions, computed by [37]:
$\langle\Phi^{N}\rangle$ $=\int_{\varphi(t=-\infty)}^{\varphi(t=\infty)}\mathcal{D}\varphi\>\prod_{i}^{N}\Phi(0,\vec{x}_{i})\>|\Psi[\varphi]|^{2}.$ (2)
Since the wavefunction serves only as an intermediate object for computing correlators, it is natural to expect that the correlators themselves may display simplifications that are otherwise absent in the wavefunction.
Both the wavefunction and the correlator admit expansion in terms of Feynman graphs, examples of which are shown in figure 1. We will focus most of our analysis at this graph level. For a graph $G$, with vertex parameters $x_{v}$ and edge parameters $y_{e}$, the wavefunction contribution takes the form:
$\Psi_{G}=\int_{0}^{\infty}dt^{V}e^{-\sum_{v}x_{v}t_{v}}\prod_{e}G_{BB}(t_{e},t^{\prime}_{e};y_{e}),$ (3)
while the corresponding correlator integral is
$\langle G\rangle=\int_{-\infty}^{\infty}dt^{V}e^{-\sum_{v}x_{v}|t_{v}|}\prod_{e}G_{F}(t_{e},t^{\prime}_{e};y_{e}).$ (4)
Here the bulk-to-bulk propagator decomposes as
$G_{BB}(t,t^{\prime};y)=G_{F}(t_{e},t^{\prime}_{e};y)-H(t,t^{\prime};y),$ (5)
where we have
$G_{F}=\frac{1}{2y}e^{-y|t-t^{\prime}|}=\frac{1}{2y}\widehat{G}_{F},\quad H=\frac{1}{2y}e^{-y(t+t^{\prime})},$ (6)
with $G_{F}$ the Feynman propagator and $H$ enforcing the vanishing condition of the bulk-to-bulk propagator when either endpoint is taken to the boundary. We emphasize that equation (4) is the textbook expression for a time-ordered correlation function. When all points are space-like separated, the operators all commute, and the time-ordering is irrelevant; thus, the expression yields the in-in correlator.
Figure 1: Example graph contributions to the wavefunction coefficients at tree-level and one-loop.
At the level of individual graphs, there is growing evidence to support this philosophical hope for simplification. First, correlators exhibit only a subset of the singularities present in the wavefunction [38], drastically reducing the complexity of the relevant function space [39, 40]. Second, the ‘melonic’ loop integrands behave much like their tree-level counterparts [37]. A key structural reason for these simplifications is that the wavefunction is computed over a half-time domain, from the asymptotic past up to the boundary. In contrast, the correlator arises from integral over the full time domain  [^1] The equivalence between in-in formalism and in-out correlation functions has been discussed in [45] for flat space, and [37] for de Sitter correlators..
In this letter, we develop these observations further and present explicit computations highlighting the relative simplicity of correlators. We will keep the discussion in flat space; however, many of these simplifications carry over to a de Sitter background, given that the integrand for conformally coupled scalars is the same, and one can obtain the integrand for minimally coupled scalars by acting with differential operators on conformally coupled scalars. Additionally, we will restrict ourselves to theories with only polynomial interactions, nevertheless also in this case we expect that the simplifications generalize for theories with derivative interactions. Our main results are:
Soft limits—At tree level when taking the external energies to zero, except those in a connected sub-graph, the tree-level correlator factorizes into a product of the amplitude obtained by shrinking the sub-graph to a vertex and the correlator of the sub-graph. In addition, at any loop order, we find that taking all but one of the vertex energies to zero, the correlator reduces to the $\ell_{0}$-integrated amplitude up to a factor. Furthermore, when taking any two external energies to zero in a tree-level correlator, it’s behaviour can be recovered from a simple contour integral.
Systematic expansion—Although the factorisation properties of the wavefunction and correlator are closely related, we show that expansions around a given pole are systematically simpler for correlators. In particular, we verify that for the correlator, the first sub-leading term always vanishes. Furthermore, for the total-energy pole, the next two subleading terms take remarkably compact forms that can be generated by acting with a simple differential operator on the amplitude.
II Wavefunction vs correlator
We work at the level of individual Feynman-diagram contributions, examples of which are shown in figure 1. For this purpose, it is convenient to introduce variables tailored to the truncated graph $G$, obtained from the original Feynman graph by deleting all external edges. To each vertex $v\in V(G)$ we assign a variable $x_{v}$, defined as the sum of the magnitude of the external momenta entering the vertex. Likewise, to each edge $e\in E(G)$ we assign a variable $y_{e}$, corresponding to the magnitude of the momentum flowing through it.
As an example, consider the five-point tree-level process shown in Fig. 1, the associated graph variables are:
$x_{1}$ $=|{\bf k}_{1}|+|{\bf k}_{2}|,\quad x_{2}=|{\bf k}_{3}|,\quad x_{3}=|{\bf k}_{4}|+|{\bf k}_{5}|,$
$y$ $=|{\bf k}_{1}+{\bf k}_{2}|=|{\bf k}_{3}+{\bf k}_{4}+{\bf k}_{5}|,$
$\tilde{y}$ $=|{\bf k}_{4}+{\bf k}_{5}|=|{\bf k}_{1}+{\bf k}_{2}+{\bf k}_{3}|.$ (7)
It will also prove useful to rescale the wavefunction and correlator as
$\Psi(G)=\frac{1}{\mathcal{N}_{G}}\widehat{\Psi}(G),\quad\langle G\rangle=\frac{1}{\mathcal{N}_{G}}\langle\widehat{G}\rangle$, (8)
where the normalization is defined as $\mathcal{N}_{G}=\prod_{e\in E(G)}2y_{e}$.
Wavefunction singularities—The wavefunction develops singularities whenever the total energy entering any connected sub-graph $\mathfrak{g}\subseteq G$ vanishes [1]. In this context, we use the convention that any edge entering the sub-graph from outside, is replaced by a half-edge attached to the corresponding vertex, with the half-edge inheriting the edge variable of the original. For instance, consider the sub-graph of the two-cycle shown in figure 1 consisting of only the top edge, for which we have:
[…]
Written in the graph variables $\{x_{v},y_{e}\}$, the singularity associated to each connected subgraph $\mathfrak{g}\subseteq G$ is:
$\mathcal{E}(\mathfrak{g})=\sum_{v\in V(\mathfrak{g})}x_{v}+\sum_{e\in H(\mathfrak{g})}y_{e},$ (10)
where $H(\mathfrak{g})$ is the set of half-edges of the subgraph. For the sub-graph considered in (9), the corresponding singularity is located at:
[…]
Wavefunction residues—Given a connected subgraph $\mathfrak{g}$ let $\{\mathfrak{h}_{i}\}$ be the connected components of the graph $G\setminus\mathfrak{g}$. For each component we define $\mathcal{Y}_{i}=H(\mathfrak{h}_{i})$ as the set of edges with one endpoint in $\mathfrak{g}$ and the other in $\mathfrak{h}_{i}$. The behaviour of the wavefunction at the pole $\mathcal{E}(\mathfrak{g})$ is as follows:
$\text{Res}_{\mathcal{E}(\mathfrak{g})=0}\widehat{\Psi}(G)=\mathcal{A}(\mathfrak{g})\prod_{i}\sum_{\sigma\in\Sigma^{|\mathcal{Y}_{i}|}}(-1)^{\sigma_{+}}\widehat{\Psi}^{\sigma}(\mathfrak{h}_{i}),$ (12)
where $\Sigma^{n}=\{+,-\}^{n}$ denotes the set of sign patterns of cardinality $n$. The shifted wavefunctions are defined as $\widehat{\Psi}^{\sigma}(\mathfrak{g})\equiv\widehat{\Psi}(\overline{\mathfrak{g}}_{i},\mathcal{Y}_{i}\rightarrow\sigma\mathcal{Y}_{i})$ where the replacement rule instructs us to flip the signs of the variables $y_{e}$ for $e\in\mathcal{Y}_{i}$ according to the sign pattern $\sigma$, and $\sigma_{+}$ denotes the number of plus signs in $\sigma$. The factor $\mathcal{A}(\mathfrak{g})$ is the amplitude of the subgraph $\mathfrak{g}$, when we are considering loop sub-graphs then it is the function obtained after integrating the zeroth component of the loop momentum ($\ell_{0}$).
Correlator singularities—The correlator contains only the singularities associated with connected, vertex-induced subgraphs of $G$, resulting in a pole structure that is considerably simpler than that of the wavefunction. This cancellation of poles mirrors the fact that the integrated wavefunction appears to have a more complicated function space than the correlator [39, 40].
As an example, the singularities for the bubble integrand are given by
[…]
This absence of poles is seen directly from the integral representation (4). It should be clear that any $2y_{e}$ appearing in a pole can only arise because a step function $\theta(t_{j}-t_{i})$ forces the integration of a given $t_{i}$ to have a boundary at $t_{j}$. To prove that this cannot happen, we need only expand each propagator of (4) in terms of step functions. Then, we can pick the terms of the integral that isolate the step-function $\theta(t_{j}-t_{i})$. There will be two terms (out of a sum of four) that could be problematic, one is:
$\int_{-\infty}^{\infty}dt_{i}dt_{j}e^{x_{i}t_{i}}e^{-x_{j}t_{j}}e^{y(t_{i}-t_{j})}\theta(t_{i}-t_{j})\theta(-t_{i})\theta(t_{j})\,$, (14)
which is zero, due to the incompatibility of the step functions. The other part is:
$\int_{-\infty}^{\infty}dt_{i}dt_{j}e^{x_{i}t_{i}}e^{-x_{j}t_{j}}e^{y(t_{i}-t_{j})}\theta(t_{i}-t_{j})\theta(t_{i})\theta(-t_{j})\,$, (15)
which once we enforce the condition from $\theta(t_{i})$ and $\theta(-t_{j})$, the step function $\theta(t_{i}-t_{j})$ becomes automatically satisfied, and no $2y$ can appear in the exponential.
Correlator residues—At tree level, the correlator exhibits a factorisation property analogous to that of the wavefunction: taking the residue at the pole $\mathcal{E}({\mathfrak{g}})$ yields
$\text{Res}_{\mathcal{E}(\mathfrak{g})=0}\langle\widehat{G}\rangle=\mathcal{A}(\mathfrak{g})\times\prod_{i}\sum_{\sigma\in\Sigma^{|\mathcal{Y}_{i}|}}\langle\hat{\mathfrak{h}}_{i}\rangle^{\sigma},$ (16)
where the shifted correlators are defined as for the wavefunction $\langle\hat{\mathfrak{h}}_{i}\rangle^{\sigma}\equiv\langle\hat{\mathfrak{h}}_{i},\mathcal{Y}_{i}\rightarrow\sigma\mathcal{Y}_{i}\rangle$. At loop level, however, the residue structure is no longer straightforward, making the simplified expansion for the correlator presented in Section IV all the more remarkable.
Melonic simplifications—As pointed out in [37], for a graph with multiple parallel edges, the correlator can expressed in terms of a single effective edge as
[…]
This allows us to restrict our analysis of the correalator to graphs in which each pair of vertices is connected by at most one edge.
III Soft Limits
In this section we study the behaviour of the correlator upon taking a subset of vertex energies to zero. We begin at tree level by introducing a useful set of variables associated to a rooted spanning tree of $G$.
Time-integrals for rooted trees—At tree-level, we may choose an arbitrary root vertex $v_{*}$ such that, for each $v\in V(G)$, there exists a unique path $P_{*}^{v}$ connecting $v_{*}$ to $v$. It is then convenient to introduce the edge variables
$s_{e}=t_{e}^{\uparrow}-t_{e}^{\downarrow},$ (18)
where $t_{e}^{\uparrow}$ denotes the time variable associated with the endpoint of $e$ further from $v_{*}$, and $t_{e}^{\downarrow}$ the time variable associated to the endpoint of $e$ closer to $v_{*}$. The translation between the edge variables $s_{e}$ and the time variables appearing in (4) is given by
$t_{v}=t_{*}+S_{v}\quad\text{ with }\quad S_{v}=\sum_{e\in P_{*}^{v}}s_{e}.$
Making these substitutions, the integral (4) becomes
$\langle\widehat{G}\rangle=\int_{-\infty}^{\infty}dt_{*}\int_{-\infty}^{\infty}ds^{E}e^{-\sum_{v}x_{v}|t_{*}+S_{v}|}\prod_{e\in E(G)}\widehat{G}_{F}(s_{e}).$ (19)
Connected subgraphs—We now explore the case where we take all vertex variables in a graph $G$ to zero, except those in a connected subgraph $\mathfrak{g}$. Choosing $x_{v_{*}}$, where $v_{*}\in\mathfrak{g}$, as the root in (19), then the integral factorizes into an integral over the variables in $\mathfrak{g}$, and an integral over the edges in $G/\mathfrak{g}$:
$\langle\widehat{G}\rangle$ $=\int\limits_{-\infty}^{\infty}dt_{*}ds_{e^{\prime}}\prod_{e^{\prime}\in E(\mathfrak{g})}\widehat{G}_{F}(s_{e^{\prime}})\int\limits_{-\infty}^{\infty}ds_{e}\prod_{e\in E(G/\mathfrak{g})}\widehat{G}_{F}(s_{e}).$
$=\langle\hat{\mathfrak{g}}\rangle\times\mathcal{A}({G/\mathfrak{g}})\,.$ (20)
In this limit the correlator factorizes into the correlator of $\mathfrak{g}$, and the amplitude of a graph where $\mathfrak{g}$ is contracted to a single vertex, which we denote by $G/\mathfrak{g}$. A particularly interesting case is when $\mathfrak{g}$ itself is a single vertex, then the property holds for loop graphs as well, and we obtain the $\ell_{0}$-integrated amplitude (and setting all the zeroth component of four-momenta to zero):
$\langle\widehat{G}\rangle|_{x_{v_{*}}\neq 0}=\frac{2}{x_{v_{*}}}\times\mathcal{A}(G),$ (21)
For a connected graph, the $l_{0}$-integrated amplitude $\mathcal{A}(G)$ satisfies a Berends-Giele type recursion
$\mathcal{A}(G)=\sum_{v\in V(G)}\prod_{\mathfrak{g}\in\kappa(G\setminus v)}\left.\frac{\mathcal{A}(\mathfrak{g})}{B(\mathfrak{g})}\right|_{x_{v}=0},$ (22)
where the product runs over the connected components of the graph $G\setminus v$. The recursion terminates on graphs consisting of a single vertex, for which we have $A(v)=1$.
Two-point functions—Now we consider the limit where we take all but two vertex variables ($x_{*}$ and $x$) to zero, in a tree graph $G$. In this limit the integral becomes
$\langle\widehat{G}\rangle=\left(\int\limits_{-\infty}^{\infty}ds_{e^{\prime}}\prod_{e\in E(G/P_{*})}\widehat{G}_{F}(s_{e^{\prime}})\right)\times$
$\times\left(\int\limits_{-\infty}^{\infty}dt_{*}ds_{e}e^{-x_{*}|t_{*}|-x|t_{*}+S_{*}|}\prod_{e\in E(P_{*})}\widehat{G}_{F}(s_{e})\right)\,.$ (23)
The first integral is simply an amplitude and we are going to get a product over $\frac{1}{y}$. The second integral can be computed by a very simple contour integral, and the result is in fact very simple (check the appendix A for the derivation), we obtain the full result as:
$\langle\widehat{G}\rangle=\left(\prod_{e\in E(G/P_{*})}\frac{1}{y_{e}}\right)\times$
$\sum_{p_{*}\in\{x_{*},x,y_{e}\}}\mathop{\mathrm{Res}}_{p\to ip_{*}}\frac{2x_{*}x\prod_{e\in E(P_{*})}2y_{e}}{(p^{2}+x_{*}^{2})(p^{2}+x^{2})\prod_{e\in E(P_{*})}(p^{2}+y_{e}^{2})}.$ (24)
IV Expanding the correlator
We now analyze the simplifications that occur in expanding the correlator about an arbitrary pole, and present a systematic expansion in the total energy.
Tree-level $\mathcal{O}(\mathcal{E}_{G})$-expansion—We begin our analysis by examining the total-energy expansion of an arbitrary tree-level graph. To bring the integral into a form suitable for expansion we introduce the rescaled variables $x_{v}=\epsilon\bar{x}_{v}$ and $\tau=\epsilon t_{*}$ such that (19) becomes
$\langle G\rangle=\frac{1}{\epsilon}\int_{-\infty}^{\infty}d\tau\int_{-\infty}^{\infty}ds^{E}F(\epsilon)\prod_{e\in E(G)}G_{F}(s_{e}),$ (25)
where we have defined
$F(\epsilon)=e^{-\sum_{v}\bar{x}_{v}|\tau+\epsilon S_{v}|}.$ (26)
The expansion of the correlator around $\epsilon=0$ is governed by the derivatives of $F(\epsilon)$, with the first few given explicitly by
$F^{\prime}(0)$ $=-F(0)\sum_{v}\bar{x}_{v}S_{v}\text{sgn}(\tau),$
$F^{\prime\prime}(0)$ $=F(0)\Bigg\{\left(\sum_{v}\bar{x}_{v}S_{v}\right)^{2}-2\sum_{v}\bar{x}_{v}S_{v}^{2}\delta(\tau)\Bigg\}.$ (27)
For instance, the leading order term is determined by the integral
$\langle G\rangle^{(-1)}=\frac{1}{\epsilon}\int_{-\infty}^{\infty}d\tau F(0)\int_{-\infty}^{\infty}ds^{E}\prod_{e\in E(G)}G_{F}(s_{e}).$ (28)
In analogy to the discussion around (19), the above integrals can be performed to find
$\langle G\rangle^{(-1)}=\frac{2}{\sum_{v}x_{v}}\prod_{e\in G}\frac{1}{y_{e}^{2}}.$ (29)
The correlator at order $\mathcal{O}(\epsilon^{0})$ is trivial to evaluate. Since $F^{\prime}(0)$ contains a factor of $\text{sgn}(\tau)$ the integrand becomes odd in $\tau$, and the integral consequently vanishes, leading to
$\langle\widehat{G}\rangle^{(0)}=0.$ (30)
We emphasise that the vanishing at next-to-leading order in the total energy expansion is unique to the correlator, making its expansion significantly simpler than that of the wavefunction!
The relative simplification continues to the next term in the expansion, which is given by:
$\langle\widehat{G}\rangle^{(1)}=\frac{-\epsilon}{\sum_{v}x_{v}}\sum_{u\neq v}x_{u}x_{v}\int_{-\infty}^{\infty}ds^{E}\sum_{e\in P_{u}^{v}}s_{e}^{2}\prod_{vv^{\prime}\in E_{G}}\widehat{G}_{F}(s_{e}),$ (31)
where $P_{u}^{v}$ is the unique path from vertex $u$ to $v$, see appendix A for the derivation. The $s_{e}^{2}$ factor is generated by the action of a suitable differential operator on the amplitude, leading to the final form of the $\mathcal{O}(\epsilon)$ contribution to the expansion:
$\langle\widehat{G}\rangle^{(1)}=-\sum_{u\neq v}\frac{x_{u}x_{v}}{\mathcal{E}_{G}}\sum_{e\in P_{u}^{v}}\partial_{y_{e}}^{2}\prod_{e\in E(G)}\frac{2}{y_{e}},$ (32)
where the total energy is given by $\mathcal{E}_{G}=\sum_{v}x_{v}$.
Loop Correlators $\mathcal{O}(\mathcal{E}_{G})$-expansion—At loop-level, our change of variables (18) imposes relations between the $s_{e}$ variables. For example, for one-loop $n$-gon diagrams, under the support of (18), for a single variable $s_{e^{\prime}}$ we must impose the constraint:
$s_{e^{\prime}}=-\sum_{e\in\mathcal{E}/e^{\prime}}s_{e}.$ (33)
For an $n$-loop diagram $n$ variables $s_{e}$ will be constrained in terms of the others. Naturally, the equations (27) still hold true, and we recover the $\ell_{0}$-integrated amplitude in this limit for the $\mathcal{E}_{G}^{-1}$ coefficient, and the $\mathcal{E}_{G}^{0}$ coefficient also vanishes. However, for the $\mathcal{E}_{G}$ there is a slight complication. We find from (47), that the cross terms of the form $s_{e}s_{e^{\prime}}$ no longer vanish, given that the integral is no longer odd. This is simply because the argument of on of the Feynman propagators will be (33) as explained above. Therefore there is no obvious form for the differential operator acting on the amplitude which gives us the $\mathcal{E}_{G}$ term. Nevertheless, it is still possible to generate this coefficient. We first notice that, because of (33), we have a choice in how we parametrize the sum of $s_{e}$ in (47). This is simply the statement that for loop diagrams there is more than one path between two sites. Then, we can consider a one higher-loop graph where we added an edge between two sites, for this diagram we can parametrize (47) such that the path between these two sites is a single $s_{e}$, and thus we do not have the issue of cross terms not vanishing. Then, for loop graphs the differential operator is simply:
$\langle\widehat{G}\rangle^{(1)}=-\frac{1}{\sum_{v}x_{v}}\sum_{u\neq v}x_{u}x_{v}\partial_{y_{uv}}^{2}\mathcal{A}^{uv}(G)|_{y_{uv}\to 0},$ (34)
where $\mathcal{A}^{uv}(G)$ is the $\ell_{0}$-integrated amplitude of the diagram with at least one edge connecting the sites $x_{u}$ and $x_{v}$. This is obvious from the integral (25) perspective, adding a single edge corresponds to adding a single Feynman propagator, when we set the additional edge to zero after acting with the differential operator we recover (47).
$\mathcal{O}(\mathcal{E}_{G}^{2})$-expansion—It turns out that the $\mathcal{O}(\mathcal{E}_{G}^{2})$ term still exhibits a much simpler structure in the correlator than in the wavefunction. After lengthy algebra, one can show that
$\langle\widehat{G}\rangle^{(2)}=\int_{-\infty}^{\infty}ds^{E}\sum_{u\neq v}\bar{x}_{u}\bar{x}_{v}\left|\sum_{e\in P_{u}^{v}}s_{e}\right|^{3}\sum\prod_{e\in E(G)}\widehat{G}_{F}(s_{e})\,.$ (35)
Now, due to the absolute value operator in the sum of $s_{e}$, even for tree level the cross terms no longer vanish, given that the integral is not odd. However, we can still apply the same solution that we had for loops, we add an edge between any two sites and then we can lower the powers of $s_{e}$ by the action of a single differential operator. Then, the $\mathcal{O}(\mathcal{E}_{G}^{2})$ term is the following:
$\langle\widehat{G}\rangle^{(2)}=-\frac{1}{\sum_{v}x_{v}}\sum_{u\neq v}x_{u}x_{v}\partial_{y_{uv}}^{3}\mathcal{A}^{uv}(G)|_{y_{uv}\to 0}.$ (36)
Beyond $\mathcal{O}(\mathcal{E}_{G}^{2})$ the coefficients become complicated polynomials of $x_{v}$, and it is no longer clear whether we can obtain them by acting with a differential operator on an amplitude.
Partial energy pole expansion—The correlator also admits a simple expansion around the partial energy poles. For a fixed connected subgraph $\mathfrak{g}$, we rescale all variables that appear in the pole as follows:
$x_{v}\rightarrow\epsilon x_{v}\text{ for all }v\in V(\mathfrak{g}),$
$y_{e}\rightarrow\epsilon y_{e}\text{ for all }e\in H(\mathfrak{g}),$ (37)
such that the partial energy becomes $\mathcal{E}_{\mathfrak{g}}\rightarrow\epsilon\mathcal{E}_{\mathfrak{g}}$. We can then derive the expansion by performing the change of variables described in (18) solely for the edges $e\in E(\mathfrak{g})$, and one time variable. At leading order we find the usual statement for the factorization of the correlator
$\langle G\rangle^{(-1)}=\frac{\mathcal{A}(\mathfrak{g})}{\mathcal{E}_{\mathfrak{g}}}\prod_{\mathfrak{h}\in\kappa(G\setminus\mathfrak{g})}\langle\mathfrak{h}\rangle,$ (38)
where the product is over all connected components of the graph $G$ with $\mathfrak{g}$ removed. Note, due to the limit we are taking, $\langle g\rangle$ is the correlator of the respective connected component with the edges connected to the central tube set to zero.
Furthermore, the first derivative with respect to $\epsilon$ is the same as in (27), therefore the sub-leading term in the series expansion around any pole of the correlator vanishes. This places strong constraints on the structure of the correlator. We have found in tree-level examples up to $4$ vertices—including both the $4$-chain and the star—that if we make an ansatz for the correlator inputing only the poles, demanding only the vanishing of subleading terms almost completely fixes the correlator.
V $\text{Tr}(\phi^{3})$ correlator
Some of the results above naturally extend to the full correlator.
Soft Limits—Recall that for a single graph, setting all vertex variables to zero except one, the correlator reduces to the corresponding contribution to the amplitude up to an overall factor. To extend this observation to the full correlator, we first recall that the vertex variables are expressed as sums over subsets of consecutive energies. Consequently, for each graph, there exists a unique vertex variable $x_{v*}$ whose sum includes the energy $|k_{*}|$. If we now switch off all other energies, this sets all vertex variables to zero except $x_{v_{*}}$ across all graphs simultaneously. The resulting expression is therefore
$\langle\phi^{n}\rangle|_{|{\bf k}_{*}|\neq 0}=\frac{1}{|{\bf k}_{*}|}A_{n},$ (39)
[…]
Pole expansion—The interesting structure in the expansion around any pole that we showed above for single graphs, generalizes for the sum of graphs. To obtain the expansion around a given pole we start by re-scaling every momenta (either external or internal), $\vec{k}_{i}\to\epsilon\,\vec{k}_{i}$, and then expand around $\epsilon=0$. The leading term will then be the usual factorization structure of the correlator, and the sub-leading term will vanish. For the total energy pole also the $\mathcal{O}(\mathcal{E}_{\text{tot}}^{1})$ term will have an interesting structure. We can reformulate our rule to obtain the operator acting on the amplitude to retrieve the $\mathcal{E}_{\text{tot}}$ term. The rule for (29) tells us that for every monomial $x_{i}\,x_{j}$ we should act with the sum over all double derivatives with respect to the every edge on the path from $x_{i}$ to $x_{j}$. To reformulate this as a full correlator rule it is best to think about it in terms of the momentum polygon for an $n$-point process. We know each triangulation of the polygon is dual to a graph, and the internal chords (parametrized by $|k_{ij}|$ of a triangulation are dual to edges (above parametrized by $y_{i}$) of the graph. Therefore, the rule of the $\mathcal{E}_{\text{tot}}$ operator is now that for each monomial $|\vec{k}_{i}|\,|\vec{k}_{j}|$ we take the sum over all the double derivatives with respect to the internal cords crossed when we draw a straight line from the outer edge $|\vec{k}_{i}|$ to the outer edge $|\vec{k}_{j}|$. This allows us to re-write (29) as:
$\nabla_{i}^{j}=-|\vec{k}_{i}|\,|\vec{k}_{j}|\sum_{e\in P_{i}^{j}}\partial_{|k|_{e}}^{2},$ (40)
where now $P_{i}^{j}$ is the set of chords which are crossed when we draw a straight line from $|\vec{k}_{i}|$ to $|\vec{k}_{j}|$ in the momentum polygon. This operator now acts on the amplitude of the full process $A_{n}$.
VI Outlook
In this letter, we explained how the relative simplicity of correlators, compared to the wavefunction, arises from the basic fact that a correlator is defined as an integration over all of spacetime. In contrast, a wavefunction is specified only on half of spacetime. Interestingly, the most physical observable—the correlator, derived from the wavefunction via the Born rule—is both ‘simpler’ and more directly related to flat-space scattering amplitudes, as opposed to the wavefunction itself, which might naively seem to have been the closer cousin.
It is important to note that this simplicity is only seen for correlation functions of the fields $\phi(x)$, that all commute at different spatial points; correlators for totally general operators, and in particular operators involving time derivatives of $\dot{\phi}(x)$ or conjugate momenta $\pi(x)$, do not exhibit the same simplicity.
As we have seen this simple picture exposes a number of remarkable properties of correlators that are obscured in the more traditional (but both less transparent and more complicated!) in-in formalism. Our investigations suggest many avenues for further research, and we close by mentioning some of the most obvious of them.
One of the most striking features of correlators we have discovered is the vanishing of subleading terms in the Laurent expansion around every pole. This turns out to be an astonishingly strong constraint in directly bootsrapping the correlator. We plan to return to a systematic analysis of the highly non-trivial power of these constraints in future work.
Two other obvious avenues of generalization are from flat-space to full cosmological correlators, as well as to derivative interactions and spinning correlators. The fundamental reason for correlator simplicity–coming from ‘full spacetime integrals’–remains in all these settings, so it will be interesting to understand the cosmological avatar of e.g. the expansion around poles we have seen in this letter.
Finally, the improved behavior of correlators upon expansion around the $\mathcal{E}_{\text{tot}}$ pole has an interesting implication for correlators in Tr $\phi^{3}$ theory. If instead of sending the total energy to zero, we shift all the internal chords of the momentum polygon to infinity, then the leading behavior at infinity also gives the amplitude, with vanishing subleading correction. Over the past few years, it has further become evident that there are hidden symmetries (“projective invariance”) that control the behavior of Tr $\phi^{3}$ amplitudes when the kinematics are shifted to infinity in special directions ($g$-vector shifts). The naive leading behavior is killed, and the subleading behavior reveals a “world at infinity”, given by amplitudes of other theories [42, 43, 44, 28]. If we instead begin not with the Tr $\phi^{3}$ amplitudes, but the correlator, then both the leading and the subleading terms vanish, and it will be fascinating to understand what the analog of the “world at infinity” is in this case.
Acknowledgements—We thank Carolina Figueiredo, Daniel Baumann and Guilherme Pimentel for useful discussions. The work of N.A.H. is supported by the DOE (Grant No. DE-SC0009988), the Simons Collaboration on Celestial Holography, the ERC UNIVERSE+ synergy grant, and the Carl B. Feinberg cross-disciplinary program in innovation at the IAS. The work of F.V. is supported by the Jonathan M. Nelson Center for Collaborative Research.
Appendix A Some Derivations
Two-point Function contour integral— Here we compute the second integral in (23), which is:
$\mathcal{I}=\int\limits_{-\infty}^{\infty}dt_{*}ds_{e}e^{-x_{*}|t_{*}|-x|t_{*}+S_{*}|}\prod_{e\in E(P_{*})}\widehat{G}_{F}(s_{e})$ (41)
By considering the Fourier transform:
$e^{-x|t_{*}+S_{*}|}=\int\limits_{-\infty}^{\infty}\frac{dp}{\pi}\frac{x}{p^{2}+x^{2}}e^{ip(t_{*}+S_{*})}\,.$ (42)
Then, every integral over $t_{*}$ and $s_{e}$ can be directly computed, for example:
$\int\limits_{-\infty}^{\infty}ds_{e}\,e^{-y|s_{e}|+ips_{e}}=\frac{2y}{p^{2}+y^{2}}\,$, (43)
therefore the first integral in (23) becomes:
$\mathcal{I}=\int\limits_{-\infty}^{\infty}\frac{dp}{\pi}\frac{2x_{*}x\prod_{e\in E(P_{*})}2y_{e}}{(p^{2}+x_{*}^{2})(p^{2}+x^{2})\prod_{e\in E(P_{*})}(p^{2}+y_{e}^{2})}\,.$ (44)
This integral is very easy to compute given that all of its poles live on the imaginary axis, we can simply compute the integral over the contour which encloses every pole in the positive imaginary axis (or equivalently all the ones in the negative axis). The integral over the arc clearly vanishes at $\pm i\infty$. Picking to contour over the positive imaginary axis, the result is a sum over all the residues on $_{*}, p=ixp=ix$ and $p=iy_{e}$:
$\mathcal{I}=\sum_{p_{*}\in\{x_{*},x,y_{e}\}}\mathop{\mathrm{Res}}_{p\to ip_{*}}\frac{2x_{*}x\prod_{e\in E(P_{*})}2y_{e}}{(p^{2}+x_{*}^{2})(p^{2}+x^{2})\prod_{e\in E(P_{*})}(p^{2}+y_{e}^{2})}$ (45)
$\mathcal{O}(E_{t})$-expansion term derivation—Starting from (25) the sub-sub-leading term is, explicitly:
$\langle G\rangle^{(1)}=\epsilon\int_{-\infty}^{\infty}d\tau\int_{-\infty}^{\infty}ds^{E}F(0)\Bigg\{\left(\sum_{v}\bar{x}_{v}S_{v}\right)^{2}-2\sum_{v}\bar{x}_{v}S_{v}^{2}\delta(\tau)\Bigg\}\prod_{e\in E(G)}G_{F}(s_{e}),$ (46)
performing the integral over $\tau$ we find that the delta function forces the integral to be $1$, and in the first term we obtain a pole over the sum of all vertex variables:
$\langle G\rangle^{(1)}=\epsilon\frac{1}{\sum_{v}\bar{x}_{v}}\int_{-\infty}^{\infty}ds^{E}\left\{\left(\sum_{v}\bar{x}_{v}S_{v}\right)^{2}-\sum_{u}\bar{x}_{u}\sum_{v}\bar{x}_{v}S_{v}^{2}\right\}\prod_{e\in E(G)}G_{F}(s_{e}).$ (47)
Recall that $S_{v}=\sum_{e\in P_{*}^{v}}s_{e}$, and $P_{*}^{v}$ is the set of edges that connect the vertex $v$ to the root vertex. It is clear that all cross terms in the square (like $s_{e}s_{e^{\prime}}$) yield odd integrands, which vanish when integrated. Then, it’s important to consider that for a given edge, $e$, we can split the graph in two connected sub-trees. One will have the root vertex, and the other will not. Then, for the edge $e$, will call the set of vertices that are in the connected sub-tree which does not contain the root vertex to be $V_{*}^{e}$, and $\bar{V}_{*}^{e}$ to be it’s complement (the vertices on the sub-tree with the root vertex). Now we can revert the sums over vertex variables, $x_{v}$, and integration variables $s_{e}$, to yield:
$\sum_{e\in E}s_{e}^{2}\left(\sum_{j\in V_{*}^{e}}\bar{x}_{j}\right)^{2}-\left(\sum_{i}^{V}x_{i}\right)\left(\sum_{e\in E}s_{e}^{2}\sum_{j\in V_{*}^{e}}\bar{x}_{j}\right)=\sum_{e\in E}s_{e}^{2}\left(\sum_{j\in V_{*}^{e}}\bar{x}_{j}\right)^{2}-\sum_{e\in E}s_{e}^{2}\left(\left(\sum_{j\in V_{*}^{e}}\bar{x}_{j}\right)^{2}+\left(\sum_{j\in V_{*}^{e}}\bar{x}_{j}\right)\left(\sum_{j\in\bar{V}_{*}^{e}}\bar{x}_{j}\right)\right)\,.$ (48)
There are three terms on the RHS, but the first and second cancel and we are left with:
$\sum_{e\in E}s_{e}^{2}\left(\sum_{u\in V_{*}^{e}}\sum_{v\in\bar{V}_{*}^{e}}\bar{x}_{u}\bar{x}_{v}\right)=-\sum_{u\neq v}x_{u}x_{v}\left(\sum_{e\in P_{u}^{v}}s_{e}^{2}\right)\,$, (49)
where we have introduced the notation $P_{u}^{v}$ for the path between $u$ to $v$. The jump from the LHS to the RHS can be seen by considering which edge integration variables $s_{e}$ are multiplying a given term $x_{u}x_{v}$. If we go back to consider the two sub-trees in which the graph splits with respect to the edge $e$, then a given $s_{e}$ in the sum can only multiply to vertex variables if the respective vertices sit on opposite sub-trees. Thus, only edges on the path from $u$ to $v$ can enter the sum on the RHS. So we are finally left with
$\langle\widehat{G}\rangle^{(1)}=\frac{-\epsilon}{\sum_{v}x_{v}}\sum_{u\neq v}x_{u}x_{v}\int_{-\infty}^{\infty}ds^{E}\sum_{e\in P_{u}^{v}}s_{e}^{2}\prod_{vv^{\prime}\in E_{G}}\widehat{G}_{F}(s_{e}),$ (50)